OpenStudy (anonymous):

simplify x^2-49/x^2+7x

hartnn (hartnn):

can u factor, x^2-49 ? 49=7^2 and \(a^2-b^2= (a+b)(a-b)\)

OpenStudy (anonymous):

so like (x-49)(x-49)

OpenStudy (anonymous):

right?

hartnn (hartnn):

not actually, it was x^2-7^2 so, (x+7)(x-7) what about x^2+7x ? factor out x...

OpenStudy (anonymous):

ok so i would do the foil method on (x+7)(x-7)

hartnn (hartnn):

no...we factored x^2-49 to (x+7)(x-7), keep it that way if you foil you will again get x^2-49... now factor x^2+7x can you ?

OpenStudy (anonymous):

no :(

OpenStudy (anonymous):

can you help me with that?

hartnn (hartnn):

just factor out the x! \(x^2+7x = x \times x + 7 \times x = x \times (x+7)\) got this ?

OpenStudy (anonymous):

oohhhhhh ok i see

hartnn (hartnn):

good, now numerator = (x+7)(x-7) and denominator = x (x+7) what gets cancelled ?

OpenStudy (anonymous):

(x+7)

hartnn (hartnn):

correct, what remains ?

OpenStudy (anonymous):

(x-7)

OpenStudy (anonymous):

so would the answer be (x-7)/x?

hartnn (hartnn):

yes! good :)

OpenStudy (anonymous):

thank you so much for your help, could you help me out with one more question?

hartnn (hartnn):

welcome ! yes, but please post it as a new question...

OpenStudy (anonymous):

ok